# Thompson sampling MAB multi-armed bandit

### v1

@@ V-A Performance for UCB-VV
In case of $\delta_{i}=0.1$, we observe the regret for $\epsilon=0.1$ is larger than for $\epsilon=0.2$. This is because a larger $\epsilon$ implies larger exploration and thus the algorithm is able to pull each arm a greater number of times and can gather more data about all the arms, before converging on the optimal arm. But excessive exploration can lead to higher regret, as we can see in the case of $\delta_{i}=0.5$. We also observe that VTS is outperforming our proposed algorithm for both sub-optimality gaps. While our proposed algorithm is optimal and offers theoretical guarantees, the superior performance of VTS highlights the need for a theoretical investigation of VTS. Thus providing the direction for future research, which could further enhance the applicability and impact of VTS. KL-UCB uses KL divergence to build confidence intervals, offering more tight bounds than traditional UCB methods that depend on simpler concentration inequalities, such as Hoeffding’s or McDiarmid’s inequality. KL-UCB is asymptotically optimal for Bernoulli bandits. This also reflects in the Fig. 1c as we compare UCB-VV with KL-UCB for $\delta=0.1$ and $\delta=0.25$ for Bernoulli distributions.
@@ V-B Performance for SHVV
In Fig. 2, we see the performance of SHVV for a fixed budget of $n=5000$ for 6 setups. Each setup presents a unique scenario, as the sub-optimality gaps follow geometric progression and arithmetic progression or can be equal within a group of sub-optimal arms. We maintain the $\sigma_{1}^{2}=\frac{1}{12}$ as our optimal arm for all cases.
@@
We observe that as the number of arms increases, $e_{n}$ increases. We see in Fig. 2a2b, 2c, 2d and 2f, an error corresponding to the detection of optimal arm increases as $K$ jumps from $16-32$ and $32-64$ for a fixed budget of $32e$ 3. The trend is similar for uniform sampling too, with SHVV having the lowest error. This implies that as arms grow, the budget required is greater than provided $n$, for the algorithm to perform optimally with growing arms.
@@ VI Case study: call option trading
Options are a widely used financial tool that investors employ to speculate on market movements or hedge against potential risks. The potential profits for option buyers are limitless, but their losses are capped, which is why they pay a premium. Conversely, option writers face potentially unlimited losses, while their profits are restricted, which is why they receive a premium. Call options are a fundamental derivative instrument in finance, granting the holder the right, but not the obligation, to purchase an underlying asset at the strike price, within a specified time period [9]. The buyer expects stock prices to rise, while the writer anticipates a decline. Volatility is a key driver of the option’s premium, as it reflects the market’s expectation of the asset’s future price fluctuations. The profit for the buyer is calculated as $\max(0,S_{t}-R)-P_{p}$, where $S_{t}$ is the stock price of the underlying asset at time $t$, $R$ is the strike price, i.e., the predetermined price at which the buyer can purchase the asset, and $P_{p}$ is the premium paid as a cost of purchasing the call option.
@@
We implement SHVV algorithm to shortlist “top 8” stocks, referring to the $8$ stocks with the highest variances given a budget of $t_{1}=20e3$. We then employ UCB-VV to select a stock for trading at time $t$ with an additional budget of $t_{2}=1e3$. If the condition $S_{t}>R$ is met, we collect the corresponding profits. This process continues iteratively: at each time step, we select the next stock according to the UCB-VV index, execute the trade if the condition holds, and accumulate the profits across all selected stocks until time horizon $n=t_{1}+t_{2}$ as shown in Table I.
@@
Figure 4: Cumulative reward (profit) v/s trading period of GBM simulated $100$ stocks for the proposed strategy v/s UCB.
@@ VII-A Proof of Theorem 1
We need to bound each $s_{i}(n)$ to bound the regret. Let $C_{t,s_{*}}=\sqrt{\frac{2\log{t}}{s_{*}}}$ be upper confidence bound. Let $z$ be any arbitrary positive integer and assume that the arm $i$ has been pulled $z$ times, including the pilot fraction. We use a subscript * to any quantity that refers to the optimal arm. At the start of the experiment, each arm is picked at least once since $S_{\rm P}=K/n$, including the sub-optimal arms. Thus,
@@
In step (a), we consider the sub-optimal arm $i$ has been picked at least $z$ times. In step (b), to pick the sub-optimal arm at time $t$, its index must be at least equal to or greater than that of the optimal arm. Step (c) involves looking back at all the optimal values we had until now, then taking their minimum, and all sub-optimal values we had, then taking their maximum. This will lead to an overcount of (b). In step (d), we upper bound by summing over all indices for the optimal and sub-optimal arms, which will include the event in step (c). Now $\bar{V}_{*}(s_{*})+C_{t,s_{*}}\leq\bar{V}_{i}(s_{i})+C_{t,s_{i}}$ implies at least one of the following must be true
@@
Similarly for $C_{t,s_{i}}$, probability of (7) will be $\leq t^{-4}$. We see that the condition in equation (8) arises when the true variances of the optimal arm and sub-optimal arm are close to each other. If the true variance of optimal is less than the true variance of any sub-optimal arm with the addition of 2 times the uncertainty term, i.e., $C_{t,s_{i}(q)}$, the agent is bound to make an error until enough number of pulls have taken place. As $z$ represents the initial number of trials, and if these trials are sufficient, then equation (8) becomes false. We see that for $z=\Big{\lceil}\frac{8\log{t}}{\delta^{2}_{i}}\Big{\rceil}$, i.e., $s_{i}=z$, the number of trials of the sub-optimal arm is substantial enough to make the occurrence of an error event to become highly unlikely. This can be seen as,
@@
$\mathcal{R}_{VV}(n)\leq 8\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\frac{\log{n}}{\delta_{i}}+\left(1+\frac{\pi^{2}}{3}\right)\left(\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\delta_{i}\right).$
@@ VII-B Proof of Lemma 2
such that $\left(\zeta>\tilde{\sigma}_{i}^{2}-\sigma^{2}_{*}>0\right)$ and $\left|I(f_{i},f_{*})-I(f_{i},\tilde{f}_{i})\right|\leq\epsilon$ for arbitrary small $\epsilon$ using assumption 1. Let $\gamma$ be the log-likelihood ratio between the $\mathcal{F}$ and $\mathcal{F}^{i}\colon\gamma=\log\frac{f_{i}\left(X_{i}\left(t_{i}(1)\right)\right)\dots f_{i}\left(X_{i}\left(t_{i}(s_{i})\right)\right)}{\tilde{f}_{i}\left(X_{i}\left(t_{i}(1)\right)\right)\dots\tilde{f}_{i}\left(X_{i}\left(t_{i}(s_{i})\right)\right)}$ where $t_{i}(s_{i})$ is the time step when the $s_{i}$-th observation of arm $i$ occurred. We show that it is unlikely to have $s_{i}<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}$ under two different scenarios for $\gamma$, i.e, $\gamma>C_{3}\log n$ and $\gamma<C_{3}\log n$.
@@
Also, using Chernoff bound $\mathbb{P}_{\mathcal{F}}[s_{i}<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})},\gamma>C_{3}\log n]$ is upper bounded for finite $n$ as
@@
$\leq\mathbb{P}_{\mathcal{F}}\bigg{[}\max_{t<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}},\sum_{s=1}^{t}\log\frac{f_{i}(X_{i}(s))}{\tilde{f}_{i}(X_{i}(s))}>C_{3}\log n\bigg{]}$
@@
$\leq\sum_{t=1}^{\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}}\exp(\frac{-a_{1}(C_{3}-C_{1})^{2})\log^{2}n}{t})$ (11)
$\leq\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}n^{-aI(f_{i},\tilde{f}_{i})\frac{(C_{3}-C_{1})^{2}}{C_{1}}}$ (12)
@@ VII-E Proof of Theorem 4
Lemma 6.
It holds that $\mathbb{P}_{k}(\xi)\geq\frac{5}{6}$.
@@
$\hskip 85.35826pt2\exp\left(\frac{-2{t}C_{4}^{2}\frac{\log(6nK)}{t}}{M^{2}}\right)$
$=2\exp\left(\frac{-2C_{4}^{2}\log(6nK)}{M^{2}}\right)$
@@
Step 2: Change of Measure
@@
Step 3: Lower Bound on $\mathbb{P}_{1}[E_{k}]$.
Assume that for the algorithm that we consider
@@
Step 4: Final bound

### v2

@@ V-A Performance for UCB-VV
In case of $\delta_{i}=0.1$, we observe the regret for $\epsilon=0.1$ is larger than for $\epsilon=0.2$. This is because a larger $\epsilon$ implies larger exploration and thus the algorithm is able to pull each arm a greater number of times and can gather more data about all the arms, before converging on the optimal arm. But excessive exploration can lead to higher regret, as we can see in the case of $\delta_{i}=0.5$. We also observe that VTS, which is a variance-based Thompson sampling (TS) algorithm, outperforms our proposed algorithm for both sub-optimality gaps. While our proposed algorithm is optimal and offers theoretical guarantees, the superior performance of VTS highlights the need for a theoretical investigation of VTS. Thus providing the direction for future research, which could further enhance the applicability and impact of VTS.
KL-UCB employs KL divergence to construct tighter confidence intervals compared to traditional UCB methods, which rely on looser concentration inequalities such as Hoeffding’s or McDiarmid’s. It is asymptotically optimal for Bernoulli bandits, which also reflects in Figure 1c. To enable a fair comparison between UCB-VV and KL-UCB, we consider a two-armed ($K=2$) Bernoulli bandit setting. Each arm is defined by its success probability $p_{1}$ and $p_{2}$, respectively. The experiment ensures both algorithms face the same sub-optimal gap $\delta$ in expected reward. For UCB-VV, setting $p_{1}=0.5$ and $p_{2}=0.1838$ yields $\delta=0.1$, accounting for variance reward. For KL-UCB, the same gap is achieved with $p_{1}=0.5$ and $p_{2}=0.4$, as it focuses solely on mean reward. In both cases, the first arm is optimal. While KL-UCB selects arms based on maximum mean reward, UCB-VV selects based on maximum variance. This distinction is illustrated in Figure 1c, where both algorithms are evaluated under $\delta=0.1$ and $\delta=0.2$.
@@ V-B Performance for SHVV
In Fig. 2, we see the performance of SHVV for a fixed budget of $n$ for 6 setups. Each setup presents a unique scenario, as the sub-optimality gaps follow geometric progression and arithmetic progression or can be equal within a group of sub-optimal arms. We maintain the $\sigma_{1}^{2}=\frac{1}{12}$ as our optimal arm for all cases.
@@
We observe that as the number of arms increases, $e_{n}$ increases. We see in Fig. 2a2b, 2c, 2d and 2f, an error corresponding to the detection of optimal arm increases as $K$ jumps from $16-32$ and $32-64$ for a fixed budget of $2000$. The trend is similar for uniform sampling, although SHVV has the lowest error. This implies that as arms grow, the budget required is greater than provided $n$, for the algorithm to perform optimally with growing arms.
@@ VI Case study: call option trading
Options are a widely used financial tool that investors employ to speculate on market movements or hedge against potential risks. The potential profits for option buyers are limitless, but their losses are capped, which is why they pay a premium. Conversely, option writers face potentially unlimited losses, while their profits are restricted, which is why they receive a premium. Call options are a fundamental derivative instrument in finance, granting the holder the right, but not the obligation, to purchase an underlying asset at the strike price, within a specified time period [9]. The buyer expects stock prices to rise, while the writer anticipates a decline. Volatility is a key driver of the option’s premium, as it reflects the market’s expectation of the asset’s future price fluctuations. The profit for the buyer is calculated as $\max(0,S_{t}-R)-P$, where $S_{t}$ is the stock price of the underlying asset at time $t$, $R$ is the strike price, i.e., the predetermined price at which the buyer can purchase the asset, and $P$ is the premium paid as a cost of purchasing the call option.
@@
We implement SHVV algorithm to shortlist “top 8” stocks referring to the $8$ stocks with the highest variances given a budget of $n_{1}=20e3$, thus pulling each arm $1674$ times. We then employ UCB-VV to select a stock for trading at time $t$ with an additional budget of $n_{2}=1000$. If the condition $S_{t}>R$ is met, we collect the corresponding profits. This process continues iteratively: at each time step, we select the next stock according to the UCB-VV index, execute the trade if the condition holds, and accumulate the profits across all selected stocks until time horizon $n=n_{1}+n_{2}$ as shown in Table I.
@@
Figure 4: Cumulative reward (profit) v/s trading period of GBM simulated $100$ stocks for the proposed strategy and UCB.
@@ VII-A Proof of Theorem 1
We need to bound each $s_{i}(n)$ to bound the regret. Let $C_{t,s_{*}}=\sqrt{\frac{2\log{t}}{s_{*}}}$ be upper confidence bound. Let $z$ be any arbitrary positive integer and assume that the arm $i$ has been pulled $z$ times, including the pilot fraction. We use a subscript * to any quantity that refers to the optimal arm. At the start of the experiment, each arm is pulled at least once since $S_{\rm P}=K/n$, including the sub-optimal arms. Thus,
@@
In step (a), we consider the sub-optimal arm $i$ has been pulled at least $z$ times. In step (b), to pull the sub-optimal arm at time $t$, its index must be at least equal to or greater than that of the optimal arm. Step (c) involves looking back at all the optimal values we had until now, then taking their minimum, and all sub-optimal values we had, then taking their maximum. This will lead to an overcount of (b). In step (d), we upper bound by summing over all indices for the optimal and sub-optimal arms, which will include the event in step (c). Now $\bar{V}_{*}(s_{*})+C_{t,s_{*}}\leq\bar{V}_{i}(s_{i})+C_{t,s_{i}}$ implies at least one of the following must be true
@@
Similarly for $C_{t,s_{i}}$, probability of (7) will be $\leq t^{-4}$. We see that the condition in equation (8) arises when the true variances of the optimal arm and sub-optimal arm are close to each other. If the true variance of optimal is less than the true variance of any sub-optimal arm with the addition of 2 times the uncertainty term, i.e., $C_{t,s_{i}(q)}$, the agent is bound to make an error until enough number of pulls have taken place. As $z$ represents the initial number of pulls, and if pulls are sufficient, then equation (8) becomes false. We see that for $z=\Big{\lceil}\frac{8\log{t}}{\delta^{2}_{i}}\Big{\rceil}$, i.e., $s_{i}=z$, the number of pulls of the sub-optimal arm is substantial enough to make the occurrence of an error event to become highly unlikely. This can be seen as,
@@
$\text{R}_{\texttt{VV}}(n)\leq 8\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\frac{\log{n}}{\delta_{i}}+\left(1+\frac{\pi^{2}}{3}\right)\left(\sum_{i:\sigma_{i}^{2}<\sigma_{*}^{2}}\delta_{i}\right).$
@@ VII-B Proof of Lemma 2
such that $\left(\zeta>\tilde{\sigma}_{i}^{2}-\sigma^{2}_{*}>0\right)$ and $\left|I(f_{i},f_{*})-I(f_{i},\tilde{f}_{i})\right|\leq\epsilon$ for arbitrary small $\epsilon$ using assumption 1. Let $\gamma$ be the log-likelihood ratio between the $\mathcal{F}$ and $\mathcal{F}^{i}\colon\gamma=\log\frac{f_{i}\left(X_{i}\left(t_{i}(1)\right)\right)\dots f_{i}\left(X_{i}\left(t_{i}(s_{i})\right)\right)}{\tilde{f}_{i}\left(X_{i}\left(t_{i}(1)\right)\right)\dots\tilde{f}_{i}\left(X_{i}\left(t_{i}(s_{i})\right)\right)}$ where $t_{i}(s_{i})$ is the time step when the $s_{i}$-th observation of arm $i$ occurred. We show that it is unlikely to have $s_{i}<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}$ under two different scenarios for $\gamma$, i.e, $\gamma>C_{3}\log n$ and $\gamma<C_{3}\log n$. Note that our proof follows a similar structure to the one presented in [37], with appropriate modifications to apply it in the context of variance.
@@
Also, using Chernoff bound, $\mathbb{P}_{\mathcal{F}}[s_{i}<\!\!\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})},\gamma>C_{3}\log n]$ is upper bounded for finite $n$ as
@@
$\leq\mathbb{P}_{\mathcal{F}}\bigg{[}\max_{t<\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}}\sum_{s=1}^{t}\log\frac{f_{i}(X_{i}(s))}{\tilde{f}_{i}(X_{i}(s))}>C_{3}\log n\bigg{]}$
@@
$\leq\sum_{t=1}^{\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}}\exp\left(\frac{-a_{1}(C_{3}-C_{1})^{2}\log^{2}n}{t}\right)$ (11)
$\leq\frac{C_{1}\log n}{I(f_{i},\tilde{f}_{i})}n^{-a_{1}I(f_{i},\tilde{f}_{i})\frac{(C_{3}-C_{1})^{2}}{C_{1}}}$ (12)
@@ VII-E Proof of Theorem 4
Step 2: It holds that $\mathbb{P}_{k}(\xi)\geq\frac{5}{6}$.
@@
$\hskip 28.45274pt2\exp\left(\frac{-2{t}C_{4}^{2}\frac{\log(6nK)}{t}}{M^{2}}\right)$
$\hskip 28.45274pt=2\exp\left(\frac{-2C_{4}^{2}\log(6nK)}{M^{2}}\right)$
@@
Step 3: Change of Measure
@@
Step 4: Lower Bound on $\mathbb{P}_{1}[E_{k}]$.
Assume that for the algorithm we consider
@@
Step 5: Final bound